# Open Communications in Nonlinear Mathematical Physics Special Issue: Bluman, 2025 ocnmp:XXXX, pp 1References

### v1

@@ Abstract
Conservation laws are computed for various nonlinear partial differential equations that arise in elasticity and acoustics. Using a scaling homogeneity approach, conservation laws are established for two models describing shear wave propagation in a circular cylinder and a cylindrical annulus. Next, using the multiplier method, conservation laws are derived for a parameterized system of constitutive equations in cylindrical coordinates involving a general expression for the Cauchy stress. Conservation laws for the Khokhlov-Zabolotskaya-Kuznetsov equation and Westervelt-type equations in various coordinate systems are also presented.
@@ 1 Introduction
Applying these methods which are implemented in the codes ConservationLawsMD.m and GeM, respectively, we compute conservation laws of various nonlinear PDEs that arise in elasticity and acoustics. In [37] we derived conservation laws based on a constitutive equation modeling stress in elastic materials with a geometry suitable for Cartesian coordinates. Here we compute conservation laws for PDEs that model the propagation of elastic waves in a circular cylinder and cylindrical annulus. More precisely, we focus on models studied by Kambapalli et al. [30] and Magan et al. [35]. These models belong to a general class of implicit constitutive equations where the strain $\epsilon$ is expressed as a non-invertible function $F(\sigma)$ of the stress $\sigma$. These so-called implicit theories originated in work of Rajagopal [43, 44] and have since been advocated by many others (see, e.g., [10, 13, 14] and the references therein).
We start with computing conservation laws for the two models reported in [30] and [35]. Both models are formulated in cylindrical coordinates and assume a power-law constitutive relation. Next, to capture other constitutive relations reported in the literature (see Table 2), we replace the power-law by an arbitrary function, $\epsilon=F(\sigma)$, relating the scalar linearized strain to Cauchy stress. The conserved densities we obtained are independent of $F$ but the fluxes depend on the integral of $F$. For completeness, we also compute conservation laws of the single wave equation for the stress that arises upon elimination of the displacement from the equations of motion.
@@
This paper is organized as follows. In Section 2, conservation laws are given for the two forementioned models for shear wave propagation in a circular cylinder and cylindrical annulus. The conservation laws for these two models are computed in Section 3 using the scaling homogeneity approach. In Section 4, conservation laws are computed with the multiplier method for a parameterized system involving an arbitrary function of stress. For specific values of the parameters, the general system includes the two previously mentioned models and also covers a broad class of models for wave propagation in elastic materials formulated in terms of cylindrical coordinates. Conservation laws for the KZK equation in Cartesian, cylindrically, and spherical coordinates are derived in Section 5. In Section 6, conservation laws are computed with the multiplier method for Westervelt-type equations in multi-space (Cartesian) coordinates as well as cylindrical and spherical coordinates. Finally, a brief discussion of the results and some conclusions are given in Section 7.
@@ 2 Model equations and some of their conservation laws
$D_{r}=\frac{\partial}{\partial r}+\sigma_{r}\frac{\partial}{\partial\sigma}+u_{r}\frac{\partial}{\partial u}+\sigma_{rr}\frac{\partial}{\partial\sigma_{r}}+u_{rr}\frac{\partial}{\partial u_{r}}+\sigma_{rt}\frac{\partial}{\partial\sigma_{t}}+u_{rt}\frac{\partial}{\partial u_{t}}+\ldots,$ (4)
@@
$D_{t}=\frac{\partial}{\partial t}+\sigma_{t}\frac{\partial}{\partial\sigma}+u_{t}\frac{\partial}{\partial u}+\sigma_{tt}\frac{\partial}{\partial\sigma_{t}}+u_{tt}\frac{\partial}{\partial u_{t}}+\sigma_{rt}\frac{\partial}{\partial\sigma_{r}}+u_{rt}\frac{\partial}{\partial u_{r}}+\ldots.$ (5)
@@
${\mathrm{D}}_{r}(r^{2}t\sigma)+{\mathrm{D}}_{t}(\delta r^{2}(u-tu_{t}))\,\dot{=}\,0,$
@@ 3 Computation of conservation laws using scaling homogeneity
In this section we discuss the scaling homogeneity approach [21, 24, 25, 37, 42] for the computation of conservation laws. To keep matters transparent we will show how to compute conservation law (8) of (1)-(2). Since that conservation law has an arbitrary $n$, we start with $n=1$ and repeat the computations for $n=2$ and $3$. Once these densities and fluxes are computed, conservation law (8) for arbitrary $n$ is obtained by pattern matching or some straightforward computation as shown at the end of this section.
@@ 3.2 Computing conservation law (8)
The key idea of the scaling homogeneity method is that all the terms in the density must have the same rank. That homogeneity in rank is because (3) should only hold on solutions of (1)-(2). Consequently, densities, fluxes, and conservation laws themselves inherit (or adopt) the scaling homogeneity of that system (and all its other continuous and discrete symmetries).
@@
Using the scaling homogeneity method and ConservationLawsMD.m, the user should select a value of $n$ (for example, $n=1$), specify the rank of the density to be computed (e.g., rank one), and the desired highest degree in $r$ and $t$ (e.g., degree 1).
@@
In other examples, the expressions for $P$ are long and complicated. Since repeated integration by parts with Mathematica might fail, ConservationLawsMD.m uses the homotopy operator [39, p. 372] to reduce the integration with respect to $t$ to a standard integral with respect to a scaling parameter $\lambda$. The homotopy operator turns out to be a very useful tool for the computation of conservation laws (see, e.g., [5, 24, 25, 26, 27, 37, 41, 42]).
@@ 4.1 Conservation laws for (1)-(2) and (10)-(11)
$\sigma\big(1+\sigma^{2}\big)^{n}$ $n\geq 0$ rational [30, [Eq. (2.17)], [31, Eq. (28)],
[33, Eqs. (4.2.7)&(5.2.6)],
[34, Eq. (2.20)], [37, Eq. (1)]
$\frac{1}{\delta}\Big(\beta\sigma+\alpha\sigma(1+\frac{\gamma}{2}\sigma^{2})^{n}\Big)$ $n$ rational [20, Eq. (2.4)], [31, Eq. (28)]
$\frac{\alpha\delta}{2}\sigma\big(1+\frac{1}{1+\sigma^{2}}\big)^{n}$ $n\geq 0$ rational [33, Eqs. (4.7.4)&(5.6.2)], [34, Eq. (2.21)],
[35, Eq. (2.9)], [46, Eq. (2.17)]
$\frac{\alpha\delta}{2}\sigma(\frac{1}{1+\sigma^{2}})^{n}$ $n\geq\frac{1}{2}$, rational [35, Eq. (2.10)]
$\frac{\alpha}{\delta}\sigma\left(\frac{1}{(1+|\sigma|^{n})^{\frac{1}{n}}}\right)$ $n>0$, integer [20, Eq. (2.5)] [46, Eq. (3.10), $n=1$]
@@
$\frac{\beta}{\delta}\sigma\big(1+\frac{1}{1+\gamma\sigma^{2}}\big)^{n}$ $n$ rational [20, Eq. (2.7)], [45, Eq. (3.13)]
$\frac{1}{\delta}{\mathcal{N}}{\mathrm{arctan}}(\theta\sigma)$ [50, Eq. (3.1)]
@@
$\frac{\alpha}{\delta}\Big(1-{\mathrm{e}}^{-\frac{\beta\sigma}{1+\tilde{\delta}|\sigma|}}\Big)$ [20, Eq. (2.6)], [45, Eq. (3.12)]
@@
$\frac{\alpha}{\delta}\Big(1-{\mathrm{e}}^{-\frac{\lambda\sigma}{1+\tilde{\delta}\sigma}}\Big)$ $n$ rational [20, Eq. (2.6), n = 1], [32, Eq. (28)],
[45, Eq. (3.14)], [47, Eq. (34), n = 1]
@@
where $\mathcal{E}_{\sigma}$ and $\mathcal{E}_{u}$ are the standard Euler operators defined in (25) and (27), respectively. Equations (44) and (45) result in an overdetermined system of PDEs for the multipliers $\Lambda^{1}$ and $\Lambda^{2}$. The derivation and solution of such determining equations has been implemented in GeM [15, 16] and other symbolic packages [53].
@@
Next, one substitutes (47) into (43) and determines $T^{r}$ and $T^{t}$. By direct integration [18], one readily obtains
@@ 6.1.1 Conservation laws of Westervelt-type equations in two dimensions
which includes the term in $\Delta_{t}\,p=\Delta p_{t}$ considered in [29, Eq. (10)] with a diffusivity-stabilizing parameter $\beta$, and an arbitrary nonlinear function $F(p)$. For completeness we kept the third-order term in time from (66) which also appears in some extended models discussed in [29, Eqs. (63) and (76)] and references therein.
@@
Equation (67) in one spatial dimension with $\alpha=0$ was studied by Márquez et al. [36, Eq. (3)]. It arises in nonlinear elasticity where $p$ then represents stress $\sigma$. See, [20, Eq. (1.1)] and [49, Eqs. (22) and (24)] with $F(p)$ as in Table 2.
@@
${\mathrm{D}}_{t}\,T^{t}+{\mathrm{D}}_{x}\,T^{x}+{\mathrm{D}}_{y}\,T^{y}=(\phi(x,y)+t\psi(x,y))\left(F(p)_{tt}-c^{2}(p_{xx}+p_{yy})\right$. (71)
$\left.-\alpha p_{ttt}-\beta(p_{xx}+p_{yy})_{t}\right),$
for arbitrary $p(x,y,t)$. By integration of (71), the density and flux can be readily derived, resulting in
@@
$\!\!\!\!\!\!\!+(pc^{2}+\beta p_{t})\left(\phi_{xx}+\phi_{yy}+t(\psi_{xx}+\psi_{yy})\right)$ (72)
@@ 6.1.3 Conservation laws of Westervelt-type equations in any number of spatial variables
Using a theorem by Igonin [28] and the direct method [4, 39], Sergyeyev [51] has derived the local conservation laws of all orders of (80) with $\alpha\neq 0$ and $\beta=0$ (i.e., the lossless case). He has also shown that for the one-dimensional Westervelt equation there are no other local conservation laws than those reported by Anco et al. [6, Eqs. (20)-(23)]. In the one-dimensional case, apart from their linearity in $t$ expressed in (82), the functions $\phi(x)$ and $\psi(x)$ are linear in $x$ because the conditions (84) reduce to $\phi_{xx}=\psi_{xx}=0$. Sergyeyev also gives a general formula for the conservation laws of all orders (see, [51, Eq. (2)]) which matches (6.1.3) again $\beta=0$ and $c=1$. Remarkably, the one-dimensional Westervelt equation has a finite number of conservation laws whereas in multiple space variables the equation has infinitely many conservation laws.
@@ 11 $r^{2}$ $r^{2}(p_{r}+\beta p_{rt})$ $r^{2}(F^{\prime}p_{t}+\alpha p_{tt})$
The conservation laws for the one-dimensional Westervelt equation in [36, Eq. (3)] are obtained by setting $m=\alpha=0$, $c=1$, F(p)=f(p)F(p)=-f(p), and replacing $r$ by $x$ in Table 6. Indeed, the first four conservation laws then correspond to those computed by Márquez et al. [36, Eqs. (21)-(24)].

### v2

# Open Communications in Nonlinear Mathematical Physics Special Issue: Bluman, 2025 ocnmp: 17124, pp 1References

@@ Abstract
Conservation laws are computed for various nonlinear partial differential equations that arise in elasticity and acoustics. Using a scaling-homogeneity approach, conservation laws are established for two models describing shear wave propagation in a circular cylinder and a cylindrical annulus. Next, using the multiplier method, conservation laws are derived for a parameterized system of constitutive equations in cylindrical coordinates involving a general expression for the Cauchy stress. Conservation laws for the Khokhlov-Zabolotskaya-Kuznetsov equation and Westervelt-type equations in various coordinate systems are also presented.
@@ 1 Introduction
Applying these methods which are implemented in the codes ConservationLawsMD.m [28] and GeM [17], respectively, we compute conservation laws of various nonlinear PDEs that arise in elasticity and acoustics. In [38] we derived conservation laws based on a constitutive equation modeling stress in elastic materials with a geometry suitable for Cartesian coordinates. Here we compute conservation laws for PDEs that model the propagation of elastic waves in a circular cylinder and cylindrical annulus. More precisely, we focus on models studied by Kambapalli et al. [31] and Magan et al. [36]. These models belong to a general class of implicit constitutive equations where the strain $\epsilon$ is expressed as a non-invertible function $F(\sigma)$ of the stress $\sigma$. These so-called implicit theories originated in work of Rajagopal [44, 45] and have since been advocated by many others (see, e.g., [10, 13, 14] and the references therein).
We start with computing conservation laws for the two models reported in [31] and [36]. Both models are formulated in cylindrical coordinates and assume a power-law constitutive relation. Next, to capture other constitutive relations reported in the literature (see Table 2), we replace the power-law by an arbitrary function, $\epsilon=F(\sigma)$, relating the scalar linearized strain to Cauchy stress. The conserved densities we obtain are independent of $F$ but the fluxes depend on the integral of $F$. For completeness, we also compute conservation laws of the single wave equation for the stress that arises upon elimination of the displacement from the equations of motion.
@@
This paper is organized as follows. In Section 2, conservation laws are given for the two forementioned models for shear wave propagation in a circular cylinder and cylindrical annulus. The conservation laws for these two models are computed in Section 3 using the scaling-homogeneity approach. In Section 4, conservation laws are computed with the multiplier method for a parameterized system involving an arbitrary function of stress covering a broad class of models for wave propagation in elastic materials formulated in terms of cylindrical coordinates. For specific values of the parameters, the general system includes the two previously mentioned models. Conservation laws for the KZK equation in Cartesian, cylindrically, and spherical coordinates are derived in Section 5. In Section 6, conservation laws are computed with the multiplier method for Westervelt-type equations in multi-space (Cartesian) coordinates as well as cylindrical and spherical coordinates. Finally, a brief discussion of the results and some conclusions are given in Section 7.
@@ 2 Model equations and some of their conservation laws
${\mathrm{D}}_{r}=\frac{\partial}{\partial r}+\sigma_{r}\frac{\partial}{\partial\sigma}+u_{r}\frac{\partial}{\partial u}+\sigma_{rr}\frac{\partial}{\partial\sigma_{r}}+u_{rr}\frac{\partial}{\partial u_{r}}+\sigma_{rt}\frac{\partial}{\partial\sigma_{t}}+u_{rt}\frac{\partial}{\partial u_{t}}+\ldots,$ (4)
@@
${\mathrm{D}}_{t}=\frac{\partial}{\partial t}+\sigma_{t}\frac{\partial}{\partial\sigma}+u_{t}\frac{\partial}{\partial u}+\sigma_{tt}\frac{\partial}{\partial\sigma_{t}}+u_{tt}\frac{\partial}{\partial u_{t}}+\sigma_{rt}\frac{\partial}{\partial\sigma_{r}}+u_{rt}\frac{\partial}{\partial u_{r}}+\ldots.$ (5)
@@
${\mathrm{D}}_{r}(r^{2}t\sigma)+{\mathrm{D}}_{t}(\delta r^{2}(u-tu_{t}))\,\dot{=}\,0,$ (7)
@@ 3 Computation of conservation laws using scaling homogeneity
In this section we discuss the scaling-homogeneity approach [21, 24, 25, 38, 43] for the computation of conservation laws. To keep matters transparent we will show how to compute conservation law (8) of (1)-(2). Since that conservation law has an arbitrary $n$, we start with $n=1$ and repeat the computations for $n=2$ and $3$. Once these densities and fluxes are computed, conservation law (8) for arbitrary $n$ is obtained by pattern matching or some straightforward computation as shown at the end of this section.
@@ 3.2 Computing conservation law (8)
The key idea of the scaling-homogeneity method is that all the terms in the density must have the same rank. That homogeneity in rank is because (3) should only hold on solutions of (1)-(2). Consequently, densities, fluxes, and conservation laws themselves inherit (or adopt) the scaling homogeneity of that system (and all its other continuous and discrete symmetries).
@@
Using the scaling-homogeneity method and ConservationLawsMD.m [28], the user should select a value of $n$ (for example, $n=1$), specify the rank of the density to be computed (e.g., rank one), and the desired highest degree in $r$ and $t$ (e.g., degree 1).
@@
In other examples, the expressions for $P$ are long and complicated. Since repeated integration by parts with Mathematica might fail, ConservationLawsMD.m uses the homotopy operator [40, p. 372] to reduce the integration with respect to $t$ to a standard integral with respect to a scaling parameter $\lambda$. The homotopy operator turns out to be a very useful tool for the computation of conservation laws (see, e.g., [5, 24, 25, 26, 27, 38, 42, 43]).
@@ 4.1 Conservation laws for (1)-(2) and (10)-(11)
$\sigma\big(1+\sigma^{2}\big)^{n}$ $n\geq 0$ rational [31, [Eq. (2.17)], [32, Eq. (28)],
[34, Eqs. (4.2.7) & (5.2.6)],
[35, Eq. (2.20)], [38, Eq. (1)]
$\frac{1}{\delta}\Big(\beta\sigma+\alpha\sigma(1+\frac{\gamma}{2}\sigma^{2})^{n}\Big)$ $n$ rational [20, Eq. (2.4)], [32, Eq. (28)]
$\frac{\alpha\delta}{2}\sigma\big(1+\frac{1}{1+\sigma^{2}}\big)^{n}$ $n\geq 0$ rational [34, Eqs. (4.7.4) & (5.6.2)], [35, Eq. (2.21)],
[36, Eq. (2.9)], [47, Eq. (2.17)]
$\frac{\alpha\delta}{2}\sigma(\frac{1}{1+\sigma^{2}})^{n}$ $n\geq\frac{1}{2}$, rational [36, Eq. (2.10)]
$\frac{\alpha}{\delta}\sigma\left(\frac{1}{(1+|\sigma|^{n})^{\frac{1}{n}}}\right)$ $n>0$, integer [20, Eq. (2.5)] [47, Eq. (3.10), $n=1$]
@@
$\frac{\beta}{\delta}\sigma\big(1+\frac{1}{1+\gamma\sigma^{2}}\big)^{n}$ $n$ rational [20, Eq. (2.7)], [46, Eq. (3.13)]
$\frac{1}{\delta}{\mathcal{N}}{\mathrm{arctan}}(\theta\sigma)$ [51, Eq. (3.1)]
@@
$\frac{\alpha}{\delta}\Big(1-{\mathrm{e}}^{-\frac{\beta\sigma}{1+\tilde{\delta}|\sigma|}}\Big)$ [20, Eq. (2.6)], [46, Eq. (3.12)]
@@
$\frac{\alpha}{\delta}\Big(1-{\mathrm{e}}^{-\frac{\lambda\sigma}{1+\tilde{\delta}\sigma}}\Big)$ $n$ rational [20, Eq. (2.6), n = 1], [33, Eq. (28)],
[46, Eq. (3.14)], [48, Eq. (34), n = 1]
@@
where $\mathcal{E}_{\sigma}$ and $\mathcal{E}_{u}$ are the standard Euler operators (see (25) and (27), respectively). Equations (44) and (45) result in an overdetermined system of PDEs for the multipliers $\Lambda^{1}$ and $\Lambda^{2}$. The derivation and solution of such determining equations has been implemented in GeM [15, 16, 17] and other symbolic packages [54].
@@
Next, one substitutes (47) into (43) and determines $T^{r}$ and $T^{t}$. Direct integration [18] as implemented in GeM [15, 16, 17] yields
@@ 6.1.1 Conservation laws of Westervelt-type equations in two dimensions
which includes the term in $\Delta_{t}\,p=\Delta p_{t}$ considered in [30, Eq. (10)] with a diffusivity-stabilizing parameter $\beta$, and an arbitrary nonlinear function $F(p)$. For completeness we kept the third-order term in time from (66) which also appears in some extended models discussed in [30, Eqs. (63) and (76)] and references therein.
@@
Equation (67) in one spatial dimension with $\alpha=0$ was studied by Márquez et al. [37, Eq. (3)]. It arises in nonlinear elasticity where $p$ then represents stress $\sigma$. See, [20, Eq. (1.1)] and [50, Eqs. (22) and (24)] with $F(p)$ as in Table 2.
@@
${\mathrm{D}}_{t}\,T^{t}+{\mathrm{D}}_{x}\,T^{x}+{\mathrm{D}}_{y}\,T^{y}=(\phi(x,y)+t\psi(x,y))\left(F(p)_{tt}-c^{2}(p_{xx}+p_{yy})\right$.
$\left.-\alpha p_{ttt}-\beta(p_{xx}+p_{yy})_{t}\right),$ (71)
for arbitrary $p(x,y,t)$. By integration of (6.1.1), the density and flux can be readily derived, resulting in
$\!\!\!\!\!\!\!{\mathrm{D}}_{t}\left[(\phi+t\psi)\left(F^{\prime}p_{t}-\alpha p_{tt}\right)-\psi(F(p)-\alpha p_{t})\right]+{\mathrm{D}}_{x}\left[-c^{2}\left[(\phi+t\psi)p_{x}-p(\phi_{x}+t\psi_{x})\right]\right$. (72)
@@
$\!\!\!\!\!\!\!+(pc^{2}+\beta p_{t})\left(\phi_{xx}+\phi_{yy}+t(\psi_{xx}+\psi_{yy})\right)$
@@ 6.1.3 Conservation laws of Westervelt-type equations in any number of spatial variables
Using a theorem by Igonin [29] and the direct method [4, 40], Sergyeyev [52] has derived the local conservation laws of all orders of (80) with $\alpha\neq 0$ and $\beta=0$ (i.e., the lossless case). He has also shown that for the one-dimensional Westervelt equation there are no other local conservation laws than those reported by Anco et al. [6, Eqs. (20)-(23)]. In the one-dimensional case, apart from their linearity in $t$ expressed in (82), the functions $\phi(x)$ and $\psi(x)$ are linear in $x$ because the conditions (84) reduce to $\phi_{xx}=\psi_{xx}=0$. Sergyeyev also gives a general formula for the conservation laws of all orders (see, [52, Eq. (2)]) which matches (6.1.3) again $\beta=0$ and $c=1$. Remarkably, the one-dimensional Westervelt equation has a finite number of conservation laws whereas in multiple space variables the equation has infinitely many conservation laws.
@@ 11 $r^{2}$ $r^{2}(p_{r}+\beta p_{rt})$ $r^{2}(F^{\prime}p_{t}+\alpha p_{tt})$
The conservation laws for the one-dimensional Westervelt equation in [37, Eq. (3)] are obtained by setting $m=\alpha=0$, $c=1$, F(p)=f(p)F(p)=-f(p), and replacing $r$ by $x$ in Table 6. Indeed, the first four conservation laws then correspond to those computed by Márquez et al. [37, Eqs. (21)-(24)].